# The Mixed Integer Trust Region Problem

### v1

@@ 2 Hardness of mixed integer trust region problems
In this short section we present some hardness results for Problems TR, MITR, and E-MIQP that we mentioned in the introduction. We recall that these three problems are listed in increasing order of generality. We start with a well-known fact about Equation TR.
Observation 1.
@@
Observation 1 also holds for the more general Problems MITR, and E-MIQP, even if $p=0$ and $n=2$. These results imply that no polynomial time algorithm, which stores numbers in binary encoding, can find an optimal solution to these problems. Next, we discuss the hardness of Equation MITR. To do so, we need a lemma on the approximation of square roots. While the square root of a rational number can be irrational, it is well-known that it can be approximated efficiently via binary search. We formalize this simple result in the next lemma. The proof is given for completeness.
@@
Let $a\in\mathbb{Q}$ be positive. There is an algorithm which, for any positive integer $k$, finds nonnegative $l,u\in\mathbb{Q}$ such that $l\leq\sqrt{a}\leq u$ and $u-l=\max\left\{1,a\right\}/2^{k}$. The running time of the algorithm is polynomial in $\operatorname{size}(a),k$.
@@
Let $l_{0}:=0$, $u_{0}:=\max\left\{1,a\right\}$. Our algorithm consists of $k$ iterations $i=1,\dots,k$. At iteration $i$, set $m_{i}:=(l_{i-1}+u_{i-1})/2$. Compare $m_{i}^{2}$ with $a$. If $m_{i}^{2}\leq a$, set $l_{i}:=m_{i}$ and $u_{i}:=u_{i-1}$. If $m_{i}^{2}>a$, set $l_{i}:=l_{i-1}$ and $u_{i}:=m_{i}$. When iteration $k$ is completed, return $l:=l_{k}$ and $u:=u_{k}$.
To show correctness of the algorithm, note that $\sqrt{a}\in[0,\max\left\{1,a\right\}]$. The algorithm performs binary search on this interval, comparing, at each iteration, the midpoint $m_{i}$ of the previous interval with $\sqrt{a}$, and updating the interval accordingly. In fact, we have $m_{i}\lesseqgtr\sqrt{a}$ if and only if $m_{i}^{2}\lesseqgtr a$. Hence, for every $i=0,1,\dots,k$, we have $\sqrt{a}\in[l_{i},u_{i}]$. Furthermore, for every $i=0,1,\dots,k$, we have $u_{i}-l_{i}=\max\left\{1,a\right\}/2^{i}$. ∎
@@
The Max-Cut problem is strongly NP-hard [10] and it is NP-hard to find an $\epsilon$-approximate solution, with $\epsilon>16/17$ [20]. It is well-known that the Max-Cut problem on a graph $G=(V,E)$ with $V=\{1,2,\dots,n\}$ and nonnegative weights $w_{uv}\in\mathbb{Z}$, for $uv\in E$, can be formulated as
@@
Observe that $\{-1,1\}^{n}=\{x\in 2\mathbb{Z}^{n}+\{1\}:x^{\mathsf{T}}x\leq r\},$ for any $r\in[n,n+8)$. We use Lemma 1 with $a:=n$ and $k:=n$. We obtain $l,u\in\mathbb{Q}$ such that $l\leq\sqrt{n}\leq u$ and $u-l=n/2^{n}$. Therefore, $u-\sqrt{n}\leq n/2^{n}<\sqrt{n+8}-\sqrt{n}$, where the last inequality can be checked with little effort. Hence we have $u\in[\sqrt{n},\sqrt{n+8})$, thus $u^{2}\in[n,n+8)$. Applying the change of variables $y:=x/u$ to the above formulation of the Max-Cut problem, we obtain the following special case of Equation MITR:
@@
In the last result of this section, we show that it is NP-hard to decide whether the feasible region of Equation MITR is nonempty. To do so, we first define the concept of lattice. Given linearly independent vectors $b^{1},\dots,b^{p}\in\mathbb{R}^{n}$, we define the lattice
@@
Then, we consider the decisional version of the closest vector problem (CVP): Given linearly independent vectors $b^{1},\dots,b^{p}\in\mathbb{Q}^{n}$, a target vector $t\in\mathbb{Q}^{n}$, and $r\in\mathbb{Q}$, determine whether there exists $y\in\Lambda(b^{1},\dots,b^{p})$ with $\lVert y-t\rVert_{2}\leq r$. This problem is well-known to be NP-hard (see theorem 3.1 in [14]). We first observe that the problem is NP-hard, even if $p=n$.
@@ 3.1 Algorithmic results
Let $x^{\dagger}$ be the feasible solution to Equation TR obtained after $k^{\prime}$ iterations with the algorithm in [24, 23]. From [23] (see in particular the first three lines of section 5), we know $f(x^{\dagger})-f(x^{\circ})\leq 2^{q-k^{\prime}}$, where $q$ is a polynomial in $\operatorname{size}(H)$, $\operatorname{size}(h)$ whose explicit form can be derived from section 4 in [23]. The result follows by setting $k^{\prime}:=\lceil k+q\rceil$, since each iteration of the algorithm performs a number of operations that is polynomial in $\operatorname{size}(H)$, $\operatorname{size}(h)$. ∎
@@
If $H$ is the zero matrix and $h$ is the zero vector, we can simply set $x^{\dagger}:=0$. Thus, in the remainder of the proof we assume that at least one among $H$ and $h$ is nonzero. We set $k:=\lceil\log(1/\epsilon)+\operatorname{size}(H)+\operatorname{size}(h)\rceil$. From Lemma 3, it suffices to show
@@
Denote by $\lambda_{\min}$ the smallest eigenvalue of $H$ and by $\lambda_{\max}$ the largest eigenvalue of $H$. Since $H$ is symmetric, we have $\lambda_{\min}=\min\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x=1\right\}$ and $\lambda_{\max}=\max\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x=1\right\}$. Since $H$ is nonzero, we have $\lVert H\rVert_{2}=\max\{|\lambda_{\min}|,\ |\lambda_{\max}|\}>0$. Every $x\in\mathbb{R}^{n}$ with $x^{\mathsf{T}}x\leq 1$ satisfies $|x^{\mathsf{T}}Hx|\leq\lVert H\rVert_{2}$, thus in the remainder of the proof it suffices to show $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. From the definition of our algorithm, it suffices to show either $|{s}^{\mathsf{T}}H{s}|\geq(1-\epsilon)\lVert H\rVert_{2}$ or $|{t}^{\mathsf{T}}H{t}|\geq(1-\epsilon)\lVert H\rVert_{2}$. In the rest of the proof, we consider separately three cases.
Note that, one of the two following cases must apply: (a) $\lVert H\rVert_{2}=|\lambda_{\min}|$ and $\lambda_{\min}<0$, or (b) $\lVert H\rVert_{2}=|\lambda_{\max}|$ and $\lambda_{\max}>0$. In fact, assume first $\lambda_{\min}<\lambda_{\max}$. If $\lVert H\rVert_{2}=|\lambda_{\min}|$, then we are in case (a). Otherwise, $\lVert H\rVert_{2}=|\lambda_{\max}|$ and we are in case (b). Assume now $\lambda_{\min}=\lambda_{\max}$. If $\lambda_{\min}<0$, then we are in case (a). Otherwise, $\lambda_{\max}>0$, and we are in case (b). Next, we consider separately cases (a) and (b).
Case (a). We have $\lambda_{\min}=\min\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x\leq 1\right\}$. From Lemma 4, we know ${s}^{\mathsf{T}}Hs-\lambda_{\min}\leq\epsilon\lVert H\rVert_{2}$. As $\lambda_{\min}=-\lVert H\rVert_{2}$, this gives $|{s}^{\mathsf{T}}H{s}|\geq-{s}^{\mathsf{T}}H{s}\geq(1-\epsilon)\lVert H\rVert_{2}$, as desired.
Case (b). We have $\lambda_{\max}=\max\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x\leq 1\right\}$. From Lemma 4, we know $-{t}^{\mathsf{T}}Ht+\lambda_{\max}\leq\epsilon\lVert H\rVert_{2}$. As $\lambda_{\max}=\lVert H\rVert_{2}$, this gives $|{t}^{\mathsf{T}}H{t}|\geq{t}^{\mathsf{T}}H{t}\geq(1-\epsilon)\lVert H\rVert_{2}$, as desired. ∎
3.2 Far points
In this section, we provide two lower bounds on the distance, in objective value, between some specific feasible points to Equation TR. First, we show that two vectors $s,t$, arbitrarily close to $h/\lVert h\rVert_{2}$ and to $-h/\lVert h\rVert_{2}$, respectively, yield $f(s)-f(t)$ arbitrarily close to $\lVert h\rVert_{2}$.
Lemma 6 (Linearly far points for Equation TR).
Let $H$ be a symmetric matrix in $\mathbb{R}^{n\times n}$, let $h\in\mathbb{R}^{n}$, and let $\epsilon\in[0,1]$. Assume $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$. Let $l\in\mathbb{R}$ be positive such that $1-\epsilon\leq l\lVert h\rVert_{2}\leq 1$. Let $s\in\mathcal{B}(lh,\epsilon)$ and $t\in\mathcal{B}(-lh,\epsilon)$. Then
$f(s)-f(t)\geq 2\left(1-4\epsilon-\epsilon^{2}\right)\lVert h\rVert_{2}.$
@@ 3.2 Far points
To show the result, we derive a lower bound for $f(s)$ and an upper bound for $f(t)$. First, we give a lower bound for $f(s)$. Let $g\in\mathbb{R}^{n}$ so that $s=lh+g$. Since $s\in\mathcal{B}(lh,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
@@
$=(lh+g)^{\mathsf{T}}H(lh+g)+h^{\mathsf{T}}(lh+g)$
$=l^{2}h^{\mathsf{T}}Hh+2lh^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+lh^{\mathsf{T}}h+h^{\mathsf{T}}g$
$\geq l^{2}h^{\mathsf{T}}Hh+2lh^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+(1-\epsilon)\lVert h\rVert_{2}+h^{\mathsf{T}}g$
$\geq l^{2}h^{\mathsf{T}}Hh-\lvert 2lh^{\mathsf{T}}Hg\rvert-\lvert g^{\mathsf{T}}Hg\rvert+(1-\epsilon)\lVert h\rVert_{2}-\lvert h^{\mathsf{T}}g\rvert$
$\geq l^{2}h^{\mathsf{T}}Hh-2l\lVert h\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}-\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}+(1-\epsilon)\lVert h\rVert_{2}-\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\geq l^{2}h^{\mathsf{T}}Hh-2\epsilon\lVert h\rVert_{2}-\epsilon^{2}\lVert h\rVert_{2}+(1-\epsilon)\lVert h\rVert_{2}-\epsilon\lVert h\rVert_{2}$
$=l^{2}h^{\mathsf{T}}Hh+\left(1-4\epsilon-\epsilon^{2}\right)\lVert h\rVert_{2}.$
Here, in the first inequality we used $l\lVert h\rVert_{2}\geq 1-\epsilon$, and in the fourth inequality $l\lVert h\rVert_{2}\leq 1$, $\lVert g\rVert_{2}\leq\epsilon$, $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$.
Similarly, we give an upper bound for $f(t)$. Redefine $g\in\mathbb{R}^{n}$ so that $t=-lh+g$. Since $t\in\mathcal{B}(-lh,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
@@
$=(-lh+g)^{\mathsf{T}}H(-lh+g)+h^{\mathsf{T}}(-lh+g)$
$=l^{2}h^{\mathsf{T}}Hh-2lh^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg-lh^{\mathsf{T}}h+h^{\mathsf{T}}g$
$\leq l^{2}h^{\mathsf{T}}Hh-2lh^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg-(1-\epsilon)\lVert h\rVert_{2}+h^{\mathsf{T}}g$
$\leq l^{2}h^{\mathsf{T}}Hh+\lvert 2lh^{\mathsf{T}}Hg\rvert+\lvert g^{\mathsf{T}}Hg\rvert-(1-\epsilon)\lVert h\rVert_{2}+\lvert h^{\mathsf{T}}g\rvert$
$\leq l^{2}h^{\mathsf{T}}Hh+2l\lVert h\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}+\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}-(1-\epsilon)\lVert h\rVert_{2}+\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\leq l^{2}h^{\mathsf{T}}Hh+2\epsilon\lVert h\rVert_{2}+\epsilon^{2}\lVert h\rVert_{2}-(1-\epsilon)\lVert h\rVert_{2}+\epsilon\lVert h\rVert_{2}$
$=l^{2}h^{\mathsf{T}}Hh-\left(1-4\epsilon-\epsilon^{2}\right)\lVert h\rVert_{2}.$
Here, in the first inequality we used $l\lVert h\rVert_{2}\geq 1-\epsilon$, and in the fourth inequality $l\lVert h\rVert_{2}\leq 1$, $\lVert g\rVert_{2}\leq\epsilon$, $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$.
@@
$f(s)-f(t)$ $\geq 2\left(1-4\epsilon-\epsilon^{2}\right)\lVert h\rVert_{2}.$
Next, we show that two vectors $u,w$, arbitrarily close to the origin, and to a vector $v$ as in Lemma 5, yield $\lvert f(w)-f(u)\rvert$ arbitrarily close to $\lVert H\rVert_{2}$.
Lemma 7 (Quadratically far points for Equation TR).
Let $H$ be a symmetric matrix in $\mathbb{R}^{n\times n}$, let $h\in\mathbb{R}^{n}$, and let $\epsilon\in[0,1]$. Assume $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Let $v\in\mathcal{B}(0,1)$ such that $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. Assume $h^{\mathsf{T}}v\geq 0$ if $v^{\mathsf{T}}Hv\geq 0$, and $h^{\mathsf{T}}v\leq 0$ if $v^{\mathsf{T}}Hv<0$. Let $w\in\mathcal{B}(v,\epsilon)$ and $u\in\mathcal{B}(0,\epsilon/2)$. Then
$\lvert f(w)-f(u)\rvert\geq\left(1-9\epsilon/2-5\epsilon^{2}/4\right)\lVert H\rVert_{2}.$
@@
$\leq\epsilon^{2}\lVert H\rVert_{2}/4+\epsilon\lVert H\rVert_{2}/2$
$=\left(\epsilon/2+\epsilon^{2}/4\right)\lVert H\rVert_{2}.$
Here, the third inequality holds because $\lVert u\rVert_{2}\leq\epsilon/2$ and $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Next, we give a lower bound for $f(w)$. Let $g\in\mathbb{R}^{n}$ so that $w=v+g$. Since $w\in\mathcal{B}(v,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.

### v2

@@ 3 Hardness of mixed integer trust region problems
In this section, we present some hardness results for Problems TR, MITR, and E-MIQP that we mentioned in the introduction. We recall that these three problems are listed in increasing order of generality. We start with a well-known fact about Equation TR.
Observation 3.
@@
Observation 3 also holds for the more general Problems MITR, and E-MIQP, even if $p=0$ and $n=2$. These results imply that no polynomial time algorithm, which stores numbers in binary encoding, can find an optimal solution to these problems. Next, we discuss the hardness of Equation MITR. To do so, we need a lemma on the approximation of square roots. While the square root of a rational number can be irrational, it is well known that it can be approximated efficiently via binary search. We formalize this simple result in the next lemma. The proof is given for completeness.
@@
Let $\alpha\in\mathbb{Q}$ be positive. There is an algorithm which, for any positive integer $k$, finds nonnegative $l,u\in\mathbb{Q}$ such that $l\leq\sqrt{\alpha}\leq u$ and $u-l=\max\left\{1,\alpha\right\}/2^{k}$. The running time of the algorithm is polynomial in $\operatorname{size}(\alpha),k$.
@@
Let $l_{0}:=0$, $u_{0}:=\max\left\{1,\alpha\right\}$. Our algorithm consists of $k$ iterations $i=1,\dots,k$. At iteration $i$, set $m_{i}:=(l_{i-1}+u_{i-1})/2$. Compare $m_{i}^{2}$ with $\alpha$. If $m_{i}^{2}\leq\alpha$, set $l_{i}:=m_{i}$ and $u_{i}:=u_{i-1}$. If $m_{i}^{2}>\alpha$, set $l_{i}:=l_{i-1}$ and $u_{i}:=m_{i}$. When iteration $k$ is completed, return $l:=l_{k}$ and $u:=u_{k}$.
To show correctness of the algorithm, note that $\sqrt{\alpha}\in[0,\max\left\{1,\alpha\right\}]$. The algorithm performs binary search on this interval, comparing, at each iteration, the midpoint $m_{i}$ of the previous interval with $\sqrt{\alpha}$, and updating the interval accordingly. In fact, we have $m_{i}\lesseqgtr\sqrt{\alpha}$ if and only if $m_{i}^{2}\lesseqgtr\alpha$. Hence, for every $i=0,1,\dots,k$, we have $\sqrt{\alpha}\in[l_{i},u_{i}]$. Furthermore, for every $i=0,1,\dots,k$, we have $u_{i}-l_{i}=\max\left\{1,\alpha\right\}/2^{i}$. ∎
@@
The Max-Cut problem is strongly NP-hard [13] and it is NP-hard to find an $\epsilon$-approximate solution, with $\epsilon>16/17$ [23]. It is well known that the Max-Cut problem on a graph $G=(V,E)$ with $V=\{1,2,\dots,n\}$ and nonnegative weights $w_{uv}\in\mathbb{Z}$, for $uv\in E$, can be formulated as
@@
Observe that $\{-1,1\}^{n}=\{x\in 2\mathbb{Z}^{n}+\{1\}:x^{\mathsf{T}}x\leq r\},$ for any $r\in[n,n+8)$. We use Lemma 1 with $\alpha:=n$ and $k:=n$. We obtain $l,u\in\mathbb{Q}$ such that $l\leq\sqrt{n}\leq u$ and $u-l=n/2^{n}$. Therefore, $u-\sqrt{n}\leq n/2^{n}<\sqrt{n+8}-\sqrt{n}$, where the last inequality can be checked with little effort. Hence we have $u\in[\sqrt{n},\sqrt{n+8})$, thus $u^{2}\in[n,n+8)$. Applying the change of variables $y:=x/u$ to the above formulation of the Max-Cut problem, we obtain the following special case of Equation MITR:
@@
Our next goal is to show that it is NP-hard to decide whether the feasible region of Equation MITR is nonempty. To do so, we first define the concept of lattice. Given linearly independent vectors $b^{1},\dots,b^{p}\in\mathbb{R}^{n}$, we define the lattice
@@
Then, we consider the decisional version of the closest vector problem (CVP): Given linearly independent vectors $b^{1},\dots,b^{p}\in\mathbb{Q}^{n}$, a target vector $t\in\mathbb{Q}^{n}$, and $r\in\mathbb{Q}$, determine whether there exists $y\in\Lambda(b^{1},\dots,b^{p})$ with $\lVert y-t\rVert_{2}\leq r$. This problem is well known to be NP-hard (see theorem 3.1 in [17]). We first observe that the problem is NP-hard, even if $p=n$.
@@
We conclude the section by observing that theorem 1 in [9] implies that the problem of deciding whether the feasible region of Equation E-MIQP is nonempty lies in the complexity class NP. Since Equation MITR can be mapped to Equation E-MIQP with the change of variables $y:=B^{-1}(x-c)$, where $B$ is the invertible matrix with columns $b^{1},\dots,b^{n}$, we obtain that also the problem of deciding whether the feasible region of Equation MITR is nonempty lies in the complexity class NP. Theorem 4 then implies that these two decision problems are NP-complete.
@@ 4.1 Algorithmic results
Let $x^{\dagger}$ be the feasible solution to Equation TR obtained after $k^{\prime}$ iterations with the algorithm in [27, 26]. From [26] (see in particular the first three lines of section 5), we know $f(x^{\dagger})-f(x^{\circ})\leq 2^{q-k^{\prime}}$, where $q$ is a polynomial in $\operatorname{size}(H)$ and $\operatorname{size}(h)$, whose explicit form can be derived from section 4 in [26]. The result follows by setting $k^{\prime}:=\lceil k+q\rceil$, since each iteration of the algorithm performs a number of operations that is polynomial in $\operatorname{size}(H)$ and $\operatorname{size}(h)$. ∎
@@
If $H$ is the zero matrix and $h$ is the zero vector, we can simply set $x^{\dagger}:=0$. Thus, in the remainder of the proof we assume that at least one among $H$ and $h$ is nonzero. We set $k:=\lceil\log(1/\epsilon)+\operatorname{size}(H)+\operatorname{size}(h)\rceil$, which can be computed in polynomial time. From Lemma 3, it suffices to show
@@
Denote by $\lambda_{\min}$ the smallest eigenvalue of $H$ and by $\lambda_{\max}$ the largest eigenvalue of $H$. Since $H$ is symmetric, we have $\lambda_{\min}=\min\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x=1\right\}$ and $\lambda_{\max}=\max\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x=1\right\}$. Since $H$ is nonzero, we have $\lVert H\rVert_{2}=\max\{|\lambda_{\min}|,\ |\lambda_{\max}|\}>0$. Every $x\in\mathbb{R}^{n}$ with $x^{\mathsf{T}}x\leq 1$ satisfies $|x^{\mathsf{T}}Hx|\leq\lVert H\rVert_{2}$, thus in the remainder of the proof it suffices to show $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. From the definition of our algorithm, it suffices to show either $|{s}^{\mathsf{T}}H{s}|\geq(1-\epsilon)\lVert H\rVert_{2}$ or $|{t}^{\mathsf{T}}H{t}|\geq(1-\epsilon)\lVert H\rVert_{2}$. In the rest of the proof, we consider separately two cases.
In the first case, we assume $\lVert H\rVert_{2}=|\lambda_{\min}|$. We then have $\lambda_{\min}<0$. Hence, $\lambda_{\min}=\min\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x\leq 1\right\}$. From Lemma 4, we know ${s}^{\mathsf{T}}Hs-\lambda_{\min}\leq\epsilon\lVert H\rVert_{2}$. As $\lambda_{\min}=-\lVert H\rVert_{2}$, this gives $|{s}^{\mathsf{T}}H{s}|\geq-{s}^{\mathsf{T}}H{s}\geq(1-\epsilon)\lVert H\rVert_{2}$, as desired.
Next, we consider the case $\lVert H\rVert_{2}=|\lambda_{\max}|$. We then have $\lambda_{\max}>0$. Hence, $\lambda_{\max}=\max\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x\leq 1\right\}$. From Lemma 4, we know $-{t}^{\mathsf{T}}Ht+\lambda_{\max}\leq\epsilon\lVert H\rVert_{2}$. As $\lambda_{\max}=\lVert H\rVert_{2}$, this gives $|{t}^{\mathsf{T}}H{t}|\geq{t}^{\mathsf{T}}H{t}\geq(1-\epsilon)\lVert H\rVert_{2}$, as desired. ∎
4.2 Feasible points to Equation TR with distant objective values
In this section, we provide two lower bounds on the difference, in objective value, between some specific feasible points to Equation TR. First, we show that two vectors $s,t$, arbitrarily close to $h/\lVert h\rVert_{2}$ and to $-h/\lVert h\rVert_{2}$, respectively, yield a lower bound on $f(s)-f(t)$ that tends to $2\lVert h\rVert_{2}$, as $\epsilon$ goes to zero. We remark that the bounds that we present are of interest when $\epsilon$ is close to zero.
Lemma 6.
Let $H$ be a symmetric matrix in $\mathbb{R}^{n\times n}$, let $h\in\mathbb{R}^{n}$, and let $\epsilon\in[0,1/2]$. Assume $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$. Let $l\in\mathbb{R}$ be positive such that $1-\epsilon\leq l\lVert h\rVert_{2}\leq 1$. Let $s\in\mathcal{B}((1-\epsilon)lh,\epsilon)$ and $t\in\mathcal{B}(-(1-\epsilon)lh,\epsilon)$. Then
$f(s)-f(t)\geq 2\left(1-5\epsilon\right)\lVert h\rVert_{2}.$
@@ 4.2 Feasible points to Equation TR with distant objective values
To show the result, we derive a lower bound for $f(s)$ and an upper bound for $f(t)$. First, we give a lower bound for $f(s)$. Let $g\in\mathbb{R}^{n}$ so that $s=(1-\epsilon)lh+g$. Since $s\in\mathcal{B}((1-\epsilon)lh,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
@@
$=((1-\epsilon)lh+g)^{\mathsf{T}}H((1-\epsilon)lh+g)+h^{\mathsf{T}}((1-\epsilon)lh+g)$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+g^{\mathsf{T}}Hg+2(1-\epsilon)lh^{\mathsf{T}}Hg+(1-\epsilon)l\lVert h\rVert_{2}^{2}+h^{\mathsf{T}}g$
$\geq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\lvert g^{\mathsf{T}}Hg\rvert-\lvert 2(1-\epsilon)lh^{\mathsf{T}}Hg\rvert+(1-\epsilon)l\lVert h\rVert_{2}^{2}-\lvert h^{\mathsf{T}}g\rvert$
$\geq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}-2(1-\epsilon)l\lVert h\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}+(1-\epsilon)l\lVert h\rVert_{2}^{2}-\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\geq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\epsilon^{2}\lVert h\rVert_{2}-2\epsilon(1-\epsilon)l\lVert h\rVert_{2}^{2}+(1-\epsilon)l\lVert h\rVert_{2}^{2}-\epsilon\lVert h\rVert_{2}$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\epsilon^{2}\lVert h\rVert_{2}+(1-2\epsilon)(1-\epsilon)l\lVert h\rVert_{2}^{2}-\epsilon\lVert h\rVert_{2}$
$\geq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\epsilon^{2}\lVert h\rVert_{2}+(1-2\epsilon)(1-\epsilon)^{2}\lVert h\rVert_{2}-\epsilon\lVert h\rVert_{2}$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\left(1-5\epsilon+4\epsilon^{2}-2\epsilon^{3}\right)\lVert h\rVert_{2}.$
Here, in the third inequality we used $\lVert g\rVert_{2}\leq\epsilon$, $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$, and in the fourth inequality $l\lVert h\rVert_{2}\geq 1-\epsilon$, since $1-2\epsilon\geq 0$.
Similarly, we give an upper bound for $f(t)$. Redefine $g\in\mathbb{R}^{n}$ so that $t=-(1-\epsilon)lh+g$. Since $t\in\mathcal{B}(-(1-\epsilon)lh,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
@@
$=(-(1-\epsilon)lh+g)^{\mathsf{T}}H(-(1-\epsilon)lh+g)+h^{\mathsf{T}}(-(1-\epsilon)lh+g)$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+g^{\mathsf{T}}Hg-2(1-\epsilon)lh^{\mathsf{T}}Hg-(1-\epsilon)l\lVert h\rVert_{2}^{2}+h^{\mathsf{T}}g$
$\leq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\lvert g^{\mathsf{T}}Hg\rvert+\lvert 2(1-\epsilon)lh^{\mathsf{T}}Hg\rvert-(1-\epsilon)l\lVert h\rVert_{2}^{2}+\lvert h^{\mathsf{T}}g\rvert$
$\leq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}+2(1-\epsilon)l\lVert h\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}-(1-\epsilon)l\lVert h\rVert_{2}^{2}+\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\leq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\epsilon^{2}\lVert h\rVert_{2}+2\epsilon(1-\epsilon)l\lVert h\rVert_{2}^{2}-(1-\epsilon)l\lVert h\rVert_{2}^{2}+\epsilon\lVert h\rVert_{2}$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\epsilon^{2}\lVert h\rVert_{2}-(1-2\epsilon)(1-\epsilon)l\lVert h\rVert_{2}^{2}+\epsilon\lVert h\rVert_{2}$
$\leq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\epsilon^{2}\lVert h\rVert_{2}-(1-2\epsilon)(1-\epsilon)^{2}\lVert h\rVert_{2}+\epsilon\lVert h\rVert_{2}$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\left(1-5\epsilon+4\epsilon^{2}-2\epsilon^{3}\right)\lVert h\rVert_{2}.$
Here, in the third inequality we used $\lVert g\rVert_{2}\leq\epsilon$, $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$, and in the fourth inequality $l\lVert h\rVert_{2}\geq 1-\epsilon$, since $1-2\epsilon\geq 0$.
@@
$f(s)-f(t)$ $\geq 2\left(1-5\epsilon+4\epsilon^{2}-2\epsilon^{3}\right)\lVert h\rVert_{2}$
$\geq 2\left(1-5\epsilon\right)\lVert h\rVert_{2}.$
Next, we show that two vectors $u,w$, arbitrarily close to the origin, and to a vector $v$ as in Lemma 5, yield a lower bound on $\lvert f(w)-f(u)\rvert$ that tends to $\lVert H\rVert_{2}$, as $\epsilon$ goes to zero.
Lemma 7.
Let $H$ be a symmetric matrix in $\mathbb{R}^{n\times n}$, let $h\in\mathbb{R}^{n}$, and let $\epsilon\in[0,1]$. Assume $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Let $v\in\mathcal{B}(0,1)$ such that $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. Assume $h^{\mathsf{T}}v\geq 0$ if $v^{\mathsf{T}}Hv\geq 0$, and $h^{\mathsf{T}}v\leq 0$ if $v^{\mathsf{T}}Hv<0$. Let $w\in\mathcal{B}((1-\epsilon)v,\epsilon)$ and $u\in\mathcal{B}(0,\epsilon)$. Then
$\lvert f(w)-f(u)\rvert\geq\left(1-7\epsilon\right)\lVert H\rVert_{2}.$
@@
$\leq\epsilon^{2}\lVert H\rVert_{2}+\epsilon\lVert H\rVert_{2}$
$=\left(\epsilon+\epsilon^{2}\right)\lVert H\rVert_{2}.$
Here, the third inequality holds because $\lVert u\rVert_{2}\leq\epsilon$ and $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Next, we give a lower bound for $f(w)$. Let $g\in\mathbb{R}^{n}$ so that $w=(1-\epsilon)v+g$. Since $w\in\mathcal{B}((1-\epsilon)v,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.